# Non-squeezing and other global rigidity results in locally conformal symplectic geometry

### v1

@@ 1. Introduction
Lcs manifolds have been studied since the 40s by several authors, among others Lee [Lee43], Libermann [Lib55], Vaisman [Vai76, Vai85], Lichnerowicz [Lic83], Banyaga [Ban02], Otiman and Stanciu [OS17], Apostolov and Dloussky [AD15, AD18], Bazzoni and Marrero [BaMa18], Chantraine and Murphy [CM19], Eliashberg and Murphy [EM23], Bertelson and Meigniez [BeMe24], Allais and Arnaud [AArn24], Chantraine and Sackel [CS24] and Currier [Cur24, Cur25]. Most of the basic notions of symplectic topology can be generalized to the lcs case. For instance, Lagrangian submanifolds and Hamiltonian isotopies on a lcs manifold $\big(M,[(\eta,\omega)]\big)$ are defined as follows. A Lagrangian submanifold is a half-dimensional submanifold on which for some (hence any) representative $(\eta,\omega)$ the lcs form $\omega$ vanishes. A vector field $X_{t}$ is said to be Hamiltonian if for some (hence any) representative $(\eta,\omega)$ the form $\iota_{X_{t}}\omega$ is $\eta$-exact. The flow of $X_{t}$ is then said to be a lcs Hamiltonian isotopy, and its time-$1$ map a lcs Hamiltonian diffeomorphism. In particular, lcs geometry can be regarded as a generalization of symplectic geometry that allows to study Hamiltonian dynamics is a more general context (and on many more spaces, since the class of manifolds admitting a lcs structure is much bigger than the class of manifolds admitting a global symplectic form [EM23, BeMe24]). The usual local flexibility properties of symplectic and contact topology (in particular the Darboux and Weinstein theorems) have natural lcs analogues. On the other hand, the study of global rigidity phenomena, which is a central theme of research in symplectic and contact topology, is less explored in the lcs case and is the main subject of the present article.
@@
The following global rigidity result in lcs geometry has been proved by Chantraine and Murphy [CM19] (see also [Cur25]). Consider a twisted cotangent bundle $T_{\beta}^{\ast}B$ with compact base $B$, and a Lagrangian submanifold $L$ in $T_{\beta}^{\ast}B$ that is lcs Hamiltonian isotopic to the zero section. Suppose that the intersections of $L$ with the zero section are transverse. Then their number is at least equal to the rank of the Novikov homology of $B$ with respect to $\beta$. Notice that this result does not imply that $L$ must always intersect the zero section, indeed the Novikov homology can vanish (for instance if $\beta$ has no zeros). In fact, it is easy to see that if $\beta$ has no zeros then there is a lcs Hamiltonian isotopy (the standard Lee flow) that displaces the zero section. The naive analogue of the Lagrangian Arnold conjecture on usual cotangent bundles, proved by Laudenbach and Sikorav [LS85], thus does not hold in the lcs case. However, the situation is similar to what already happens in the contact case for $1$-jet bundles, where the standard Reeb flow also displaces the zero section. In the contact case the following analogue of the Lagrangian Arnold conjecture, where intersections are replaced by Reeb chords, has been proved by Chekanov [Che96] and Chaperon [Cha95]. We endow the $1$-jet bundle $J^{1}B=T^{\ast}B\times\mathbb{R}$ with its standard contact structure $\xi=\ker(\lambda-dz)$, where $z$ is the coordinate on $\mathbb{R}$. If $B$ is compact then for any Legendrian submanifold $\Lambda$ of $J^{1}B$ that is the image of the zero section by a contact isotopy the number of Reeb chords with respect to $\lambda-dz$ between $\Lambda$ and the zero section is greater than or equal to the cup-length of $B$, and greater than or equal to the sum of the Betti numbers if all the Reeb chords are transverse. Our first observation is that a similar result also holds in the lcs case for essential Lee chords on twisted cotangent bundles. Before precisely stating the result, we need to recall or introduce a few more notions, referring to Sections 2 and 3 for more details. The Lee flow on a lcs manifold $\big(M,[(\eta,\omega)]\big)$ with respect to a representative $(\eta,\omega)$ is the flow $\{\varphi_{t}^{\omega}\}$ of the vector field $R_{\omega}$ defined by the relation $\omega(R_{\omega},\,\cdot\,)=\eta$ (equivalently, the lcs Hamiltonian isotopy with Hamiltonian function $1$ with respect to $(\eta,\omega)$). For instance, the Lee flow on a locally conformal symplectization $\big(S^{1}\times Y\,,\,[(-d\theta,d_{-d\theta}\alpha)]\big)$ with respect to $(-d\theta,d_{-d\theta}\alpha)$ is $\{\operatorname{id}\times\varphi_{t}^{\alpha}\}$, where $\{\varphi_{t}^{\alpha}\}$ is the Reeb flow on $Y$ with respect to $\alpha$, and the Lee flow on a twisted cotangent bundle $T^{\ast}_{\beta}B$ with respect to $(\pi^{\ast}\beta,d_{\pi^{\ast}\beta}\lambda)$ is given by $\varphi_{t}^{\omega}(\sigma_{q})=\sigma_{q}+t\beta(q)$. Let $L_{1}$ and $L_{2}$ be two Lagrangian submanifolds of a lcs manifold $\big(M,[(\eta,\omega)]\big)$. A Lee chord with respect to $(\eta,\omega)$ from $L_{1}$ to $L_{2}$ is a pair $(T,\gamma)$ with $T$ a real number (called the time-shift of the Lee chord) and $\gamma:[0,1]\rightarrow M$ a curve such that $\gamma(0)\in L_{1}$, $\gamma(1)\in L_{2}$ and $\dot{\gamma}(t)=T\,R_{\omega}\big(\gamma(t)\big)$ for all $t$. We say that a Lee chord $(T,\gamma)$ with respect to $(\eta,\omega)$ from $L_{1}$ to $L_{2}$ is transverse if $(\varphi_{T}^{\omega})_{\ast}\big(T_{\gamma(0)}L_{1}\big)+T_{\gamma(1)}L_{2}=T_{\gamma(1)}M$. A lcs manifold $\big(M,[(\eta,\omega)]\big)$ is said to be exact if for some (hence any) representative $(\eta,\omega)$ the lcs form $\omega$ is $\eta$-exact. Any $1$-form $\lambda$ such that $\omega=d_{\eta}\lambda$ is then called a Liouville form for $(\eta,\omega)$. If $\lambda$ is a Liouville form for $(\eta,\omega)$ then, for every function $f$, $e^{f}\lambda$ is a Liouville form for $(\eta+df,e^{f}\omega)$. A Liouville lcs manifold is a manifold $M$ endowed with an equivalence class $(\eta,\omega,\lambda)]$ of triples formed by a closed $1$-form $\eta$, a non-degenerate $2$-form $\omega$ and a $1$-form $\lambda$ such that $\omega=d_{\eta}\lambda$, under the equivalence relation $(\eta,\omega,\lambda)\sim(\eta+df,e^{f}\omega,e^{f}\lambda)$. For instance, twisted cotangent bundles $\big(T^{\ast}B\,,\,[(\pi^{\ast}\beta,d_{\pi^{\ast}\beta}\lambda,\lambda)]\big)$ and locally conformal symplectizations $\big(S^{1}\times Y\,,\,[(-d\theta,d_{-d\theta}\alpha,\alpha)]\big)$ are Liouville lcs manifolds. A Lagrangian submanifold $L$ of a Liouville lcs manifold $\big(M,[(\eta,\omega,\lambda)]\big)$ is said to be exact if the $1$-form $i^{\ast}\lambda$, where $i:L\rightarrow M$ is the inclusion, is $i^{\ast}\eta$-exact. In particular, any Lagrangian submanifold of $T_{\beta}^{\ast}B$ that is the image of the zero section by a lcs Hamiltonian diffeomorphism is exact (Corollary 2.32). We say that an exact Lagrangian submanifold $L$ of a Liouville lcs manifold $\big(M,[(\eta,\omega,\lambda)]\big)$ has action $S_{L}:L\rightarrow\mathbb{R}$ with respect to $(\eta,\omega,\lambda)$ if $i^{\ast}\lambda=d_{i^{\ast}\eta}S_{L}$. If $L$ is connected and the restriction of $\eta$ to $L$ is not exact then such action $S_{L}$ is uniquely defined (see Remark 2.1). In this case we define the action with respect to $(\eta,\omega,\lambda)$ of a Lee chord $(T,\gamma)$ between two exact Lagrangian submanifolds $L_{1}$ and $L_{2}$ to be the real number
@@
    Let $\varphi$ be the time-$1$ map of a sufficiently $\mathcal{C}^{0}$-small lcs Hamiltonian isotopy of a compact connected Liouville lcs manifold $\big(M,[(\eta,\omega,\lambda)]\big)$ with $\eta$ not exact. Then the number of essential translated points of $\varphi$ with respect to any representative $(\eta,\omega)$ is greater than or equal to the cup-length of $M$, and greater than or equal to the sum of the Betti numbers of $M$ if all the essential Lee chords for the essential translated points are non-degenerate.
@@
In order to prove Theorem 1.2 (ii) and Theorem 1.3 (ii) we use generating functions to define spectral selectors for compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n+1}$ and $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, analogous to the spectral selectors for compactly supported Hamiltonian diffeomorphisms of $\mathbb{R}^{2n}$ defined by Viterbo [Vit92] and the spectral selectors for compactly supported contactomorphisms contact isotopic to the identity of $\mathbb{R}^{2n+1}$ and $\mathbb{R}^{2n}\times S^{1}$ defined respectively by Bhupal [Bhu01] and the third author [San10, San11a]. The key observation is that for the lcs manifold $S^{1}\times\mathbb{R}^{2n+1}$ we do not just have the local identifications of neighborhoods of Lagrangian submanifolds of the lcs product with neighborhoods of the zero section of a twisted cotangent bundle given by the lcs Weinstein theorem, but also a global identification of the whole lcs product with the twisted cotangent bundle $T_{-d\theta}(S^{1}\times\mathbb{R}^{2n+1})$, where $\theta$ denotes the coordinate on $S^{1}$. Indeed, the lcs product of $S^{1}\times\mathbb{R}^{2n+1}$ can be identified to the locally conformal symplectization $S^{1}\times(\mathbb{R}^{2n+1}\times\mathbb{R}^{2n+1}\times\mathbb{R})$ of the contact product of $\mathbb{R}^{2n+1}$ (Proposition 2.41), and the twisted cotangent bundle $T_{-d\theta}(S^{1}\times\mathbb{R}^{2n+1})$ can be identified to the locally conformal symplectization $S^{1}\times J^{1}\mathbb{R}^{2n+1}$ of $J^{1}\mathbb{R}^{2n+1}$ (Lemma 2.13). We can thus lift the strict contactomorphism from $\mathbb{R}^{2n+1}\times\mathbb{R}^{2n+1}\times\mathbb{R}$ to $J^{1}\mathbb{R}^{2n+1}$ used in [Bhu01, San10, San11a] to the locally conformal symplectizations to obtain the desired identification (see equation (9) in Section 5). By considering the image by this identification of the lcs graph of a lcs Hamiltonian diffeomorphism $\varphi$ of $S^{1}\times\mathbb{R}^{2n+1}$ we obtain a Lagrangian submanifold $L_{\varphi}$ of $T_{-d\theta}(S^{1}\times\mathbb{R}^{2n+1})$. If $\varphi$ is compactly supported then $L_{\varphi}$ can be compactified to a Lagrangian submanifold $\overline{L_{\varphi}}$ of $T_{-d\theta}(S^{1}\times S^{2n+1})$, lcs Hamiltonian isotopic to the zero section. Similarly, we can associate to a compactly supported lcs Hamiltonian diffeomorphism $\varphi$ of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$ a Lagrangian submanifold $L_{\varphi}$ of $T_{-d\theta}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$, which can be compactified to a Lagrangian submanifold $\overline{L_{\varphi}}$ of $T_{-d\theta}(S^{1}\times S^{2n}\times S^{1})$, lcs Hamiltonian isotopic to the zero section. Using generating functions quadratic at infinity for $\overline{L_{\varphi}}$ we define, exactly as in the contact case for contactomorphisms of $\mathbb{R}^{2n+1}$ and $\mathbb{R}^{2n}\times S^{1}$ [Bhu01, San11a, San10] spectral selectors $c_{+}(\varphi)\geq 0$ and $c_{-}(\varphi)\leq 0$ for any compactly supported lcs Hamiltonian diffeomorphism $\varphi$ of $S^{1}\times\mathbb{R}^{2n+1}$ or $S^{1}\times\mathbb{R}^{2n}\times S^{1}$. These spectral selectors have the same properties as in the contact case (Proposition 6.3), and thus can be used to obtain similar applications. To start with, spectrality ($c_{+}(\varphi)$ and $c_{-}(\varphi)$ are equal to the action of an essential translated point of $\varphi$) and non-degeneracy ($c_{+}(\varphi)=c_{-}(\varphi)=0$ if and only if $\varphi$ is the identity) imply Theorem 1.2 (ii). Moreover, as we will see, the properties listed in Proposition 6.3 imply that the partial relation $\preceq$ on the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n+1}$ or $S^{1}\times\mathbb{R}^{2n}\times S^{1}$ defined by posing $\varphi_{1}\preceq\varphi_{2}$ if $c_{+}(\varphi_{1}\circ\varphi_{2}^{-1})=0$ is a bi-invariant partial order (Theorem 7.1). By comparing the partial relation $\preceq$ with the one $\leq$ used to define orderability we then obtain a proof of Theorem 1.3 (ii). Finally, we show that the arguments in [San11a, San10] can be generalized to the lcs case to obtain an unbounded integer-valued bi-invariant metric $d$ on the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, compatible with the partial orders $\preceq$ and $\leq$, by posing
@@
(Theorems 7.3 and 7.4), and an integer-valued lcs capacity for domains, by posing
@@
This will be the object of a forthcoming work.
@@
The article is organized as follows. In Section 2 we gather the basic notions on lcs manifolds that are needed in the rest of the article. Moreover, we discuss in 2.10 the notion of orderability and prove Theorem 1.3 (i). In Section 3 we define essential Lee chords and translated points. In Section 4 we recall from [CM19] the theory of generating functions for Lagrangian submanifolds of twisted cotangent bundles, and prove Theorem 1.1 and Theorem 1.2 (i). In Section 5 we define generating functions for compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n+1}$ and $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, and in Section 6 we use them to define the spectral selectors $c_{+}$ and $c_{-}$. A first application of these spectral selectors, Theorem 1.2 (ii), is also discussed in Section 6. In Section 7 we use the spectral selectors to define a partial order on the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n+1}$ and $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, and a bi-invariant metric on the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, and obtain a proof of Theorem 1.3 (ii). In Section 8 we define a lcs capacity for domains of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, and use it to prove the lcs non-squeezing theorem for integers (Theorem 1.4). Finally, in Section 9 we discuss, following [SS24], how this non-squeezing result implies that the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$ does not have the Rokhlin property.

### v2

@@ 1. Introduction
Lcs manifolds have been studied since the 40s by several authors, among others Lee [Lee43], Libermann [Lib55], Vaisman [Vai76, Vai85], Lichnerowicz [Lic83], Banyaga [Ban02], Otiman and Stanciu [OS17], Apostolov and Dloussky [AD15, AD18], Bazzoni and Marrero [BaMa18], Chantraine and Murphy [CM19], Eliashberg and Murphy [EM23], Bertelson and Meigniez [BeMe24], Gironella and Toussaint [GT24] Allais and Arnaud [AArn24], Chantraine and Sackel [CS24], Currier [Cur24, Cur25] and Van Overschelde [VO25a, VO25b]. Most of the basic notions of symplectic topology can be generalized to the lcs case. For instance, Lagrangian submanifolds and Hamiltonian isotopies on a lcs manifold $\big(M,[(\eta,\omega)]\big)$ are defined as follows. A Lagrangian submanifold is a half-dimensional submanifold on which for some (hence any) representative $(\eta,\omega)$ the lcs form $\omega$ vanishes. A vector field $X_{t}$ is said to be Hamiltonian if for some (hence any) representative $(\eta,\omega)$ the form $\iota_{X_{t}}\omega$ is $\eta$-exact. The flow of $X_{t}$ is then said to be a lcs Hamiltonian isotopy, and its time-$1$ map a lcs Hamiltonian diffeomorphism. In particular, lcs geometry can be regarded as a generalization of symplectic geometry that allows to study Hamiltonian dynamics in a more general context (and on many more spaces, since the class of manifolds admitting a lcs structure is much bigger than the class of manifolds admitting a global symplectic form [EM23, BeMe24]). Another important reason for being interested in locally conformal symplectic structures comes from the hope that, in higher dimension, taut codimension one foliations can be deformed into tight contact structures, as in dimension 3 (cf. [ET98]). If the foliation is defined by a $1$-form $\alpha$, with holonomy $1$-form $\eta$ ($d\alpha=\eta\wedge\alpha$) the easiest way to deform the foliation is to deform $\alpha$, that is consider a smooth family of $1$-forms $\alpha_{t}$ with $\alpha_{0}=\alpha$. If the first order of this deformation $\dot{\alpha}_{t}$ is a $\eta$-Liouville form, then the deformation is a contact structure for all sufficiently small positive $t$’s. This was the motivation of the authors of [BeMe24] for their construction of (foliated) exact lcs form with a fixed (leafwise) Lee form.
The usual local flexibility properties of symplectic and contact topology (in particular, the Darboux and Weinstein theorems) have natural lcs analogues. One of the central results of symplectic geometry, $\mathcal{C}^{0}$-rigidity of symplectomorphisms, also readily implies its lcs counterpart (see Section 2.5). On the other hand, the study of other global rigidity phenomena is less explored in the lcs case. One of the reasons is the absence of a Stokes theorem for the Lichnerowicz–de Rham differential, and thus the lack of $C^{0}$-bounds on the energy of pseudo-holomorphic curves, which proves to be a major hurdle to studying deformations of holomorphic curves on these manifolds (see Section 2.5 in [CM19]).
@@
However, as noticed by Chantraine and Murphy [CM19], the theory of generating functions adapts elegantly to the lcs case. In particular, Chantraine and Murphy use generating functions to prove the following global rigidity result (see also [Cur25] for refinements). Consider a twisted cotangent bundle $T_{\beta}^{\ast}B$ with compact base $B$, and a Lagrangian submanifold $L$ in $T_{\beta}^{\ast}B$ that is lcs Hamiltonian isotopic to the zero section. Suppose that the intersections of $L$ with the zero section are transverse. Then their number is at least equal to the rank of the Novikov homology of $B$ with respect to $\beta$. Notice that this result does not imply that $L$ must always intersect the zero section, indeed the Novikov homology can vanish (for instance if $\beta$ has no zeros). In fact, it is easy to see that if $\beta$ has no zeros then there is a lcs Hamiltonian isotopy (the standard Lee flow) that displaces the zero section. The naive analogue of the Lagrangian Arnold conjecture on usual cotangent bundles, proved by Laudenbach and Sikorav [LS85], thus does not hold in the lcs case. However, the situation is similar to what already happens in the contact case for $1$-jet bundles, where the standard Reeb flow also displaces the zero section. In the contact case the following analogue of the Lagrangian Arnold conjecture, where intersections are replaced by Reeb chords, has been proved by Chekanov [Che96] and Chaperon [Cha95]. We endow the $1$-jet bundle $J^{1}B=T^{\ast}B\times\mathbb{R}$ with its standard contact structure $\xi=\ker(\lambda-dz)$, where $z$ is the coordinate on $\mathbb{R}$. If $B$ is compact then for any Legendrian submanifold $\Lambda$ of $J^{1}B$ that is the image of the zero section by a contact isotopy the number of Reeb chords with respect to $\lambda-dz$ between $\Lambda$ and the zero section is greater than or equal to the cup-length of $B$, and greater than or equal to the sum of the Betti numbers if all the Reeb chords are transverse.
Our first observation is that a similar result also holds in the lcs case for essential Lee chords on twisted cotangent bundles. Before precisely stating the result, we need to recall or introduce a few more notions, referring to Sections 2 and 3 for more details. The Lee flow on a lcs manifold $\big(M,[(\eta,\omega)]\big)$ with respect to a representative $(\eta,\omega)$ is the flow $\{\varphi_{t}^{\omega}\}$ of the vector field $R_{\omega}$ defined by the relation $\omega(R_{\omega},\,\cdot\,)=\eta$ (equivalently, the lcs Hamiltonian isotopy with Hamiltonian function $1$ with respect to $(\eta,\omega)$). For instance, the Lee flow on a locally conformal symplectization $\big(S^{1}\times Y\,,\,[(-d\theta,d_{-d\theta}\alpha)]\big)$ with respect to $(-d\theta,d_{-d\theta}\alpha)$ is $\{\operatorname{id}\times\varphi_{t}^{\alpha}\}$, where $\{\varphi_{t}^{\alpha}\}$ is the Reeb flow on $Y$ with respect to $\alpha$, and the Lee flow on a twisted cotangent bundle $T^{\ast}_{\beta}B$ with respect to $(\pi^{\ast}\beta,d_{\pi^{\ast}\beta}\lambda)$ is given by $\varphi_{t}^{\omega}(\sigma_{q})=\sigma_{q}+t\beta(q)$. Let $L_{1}$ and $L_{2}$ be two Lagrangian submanifolds of a lcs manifold $\big(M,[(\eta,\omega)]\big)$. A Lee chord with respect to $(\eta,\omega)$ from $L_{1}$ to $L_{2}$ is a pair $(T,\gamma)$ with $T$ a real number (called the time-shift of the Lee chord) and $\gamma:[0,1]\rightarrow M$ a curve such that $\gamma(0)\in L_{1}$, $\gamma(1)\in L_{2}$ and $\dot{\gamma}(t)=T\,R_{\omega}\big(\gamma(t)\big)$ for all $t$. We say that a Lee chord $(T,\gamma)$ with respect to $(\eta,\omega)$ from $L_{1}$ to $L_{2}$ is transverse if $(\varphi_{T}^{\omega})_{\ast}\big(T_{\gamma(0)}L_{1}\big)+T_{\gamma(1)}L_{2}=T_{\gamma(1)}M$. A lcs manifold $\big(M,[(\eta,\omega)]\big)$ is said to be exact if for some (hence any) representative $(\eta,\omega)$ the lcs form $\omega$ is $\eta$-exact. Any $1$-form $\lambda$ such that $\omega=d_{\eta}\lambda$ is then called a Liouville form for $(\eta,\omega)$. If $\lambda$ is a Liouville form for $(\eta,\omega)$ then, for every function $f$, $e^{f}\lambda$ is a Liouville form for $(\eta+df,e^{f}\omega)$. A Liouville lcs manifold is a manifold $M$ endowed with an equivalence class $(\eta,\omega,\lambda)]$ of triples formed by a closed $1$-form $\eta$, a non-degenerate $2$-form $\omega$ and a $1$-form $\lambda$ such that $\omega=d_{\eta}\lambda$, under the equivalence relation $(\eta,\omega,\lambda)\sim(\eta+df,e^{f}\omega,e^{f}\lambda)$. For instance, twisted cotangent bundles $\big(T^{\ast}B\,,\,[(\pi^{\ast}\beta,d_{\pi^{\ast}\beta}\lambda,\lambda)]\big)$ and locally conformal symplectizations $\big(S^{1}\times Y\,,\,[(-d\theta,d_{-d\theta}\alpha,\alpha)]\big)$ are Liouville lcs manifolds. A Lagrangian submanifold $L$ of a Liouville lcs manifold $\big(M,[(\eta,\omega,\lambda)]\big)$ is said to be exact if the $1$-form $i^{\ast}\lambda$, where $i:L\rightarrow M$ is the inclusion, is $i^{\ast}\eta$-exact. In particular, any Lagrangian submanifold of $T_{\beta}^{\ast}B$ that is the image of the zero section by a lcs Hamiltonian diffeomorphism is exact (Corollary 2.33). We say that an exact Lagrangian submanifold $L$ of a Liouville lcs manifold $\big(M,[(\eta,\omega,\lambda)]\big)$ has action $S_{L}:L\rightarrow\mathbb{R}$ with respect to $(\eta,\omega,\lambda)$ if $i^{\ast}\lambda=d_{i^{\ast}\eta}S_{L}$. If $L$ is connected and the restriction of $\eta$ to $L$ is not exact then such action $S_{L}$ is uniquely defined (see Remark 2.1). In this case we define the action with respect to $(\eta,\omega,\lambda)$ of a Lee chord $(T,\gamma)$ between two exact Lagrangian submanifolds $L_{1}$ and $L_{2}$ to be the real number
@@
    Let $\varphi$ be the time-$1$ map of a sufficiently $\mathcal{C}^{0}$-small lcs Hamiltonian isotopy of a compact connected Liouville lcs manifold $\big(M,[(\eta,\omega,\lambda)]\big)$ with $\eta$ not exact. Then the number of essential translated points of $\varphi$ with respect to any representative $(\eta,\omega,\lambda)$ is greater than or equal to the cup-length of $M$, and greater than or equal to the sum of the Betti numbers of $M$ if all the essential Lee chords for the essential translated points are non-degenerate.
@@
In order to prove Theorem 1.2 (ii) and Theorem 1.3 (ii) we use generating functions to define spectral selectors for compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n+1}$ and $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, analogous to the spectral selectors for compactly supported Hamiltonian diffeomorphisms of $\mathbb{R}^{2n}$ defined by Viterbo [Vit92] and the spectral selectors for compactly supported contactomorphisms contact isotopic to the identity of $\mathbb{R}^{2n+1}$ and $\mathbb{R}^{2n}\times S^{1}$ defined respectively by Bhupal [Bhu01] and the third author [San10, San11a]. The key observation is that for the lcs manifold $S^{1}\times\mathbb{R}^{2n+1}$ we do not just have the local identifications of neighborhoods of Lagrangian submanifolds of the lcs product with neighborhoods of the zero section of a twisted cotangent bundle given by the lcs Weinstein theorem, but also a global identification of the whole lcs product with the twisted cotangent bundle $T_{-d\theta}(S^{1}\times\mathbb{R}^{2n+1})$, where $\theta$ denotes the coordinate on $S^{1}$. Indeed, the lcs product of $S^{1}\times\mathbb{R}^{2n+1}$ can be identified to the locally conformal symplectization $S^{1}\times(\mathbb{R}^{2n+1}\times\mathbb{R}^{2n+1}\times\mathbb{R})$ of the contact product of $\mathbb{R}^{2n+1}$ (Proposition 2.42), and the twisted cotangent bundle $T_{-d\theta}(S^{1}\times\mathbb{R}^{2n+1})$ can be identified to the locally conformal symplectization $S^{1}\times J^{1}\mathbb{R}^{2n+1}$ of $J^{1}\mathbb{R}^{2n+1}$ (Lemma 2.13). We can thus lift the strict contactomorphism from $\mathbb{R}^{2n+1}\times\mathbb{R}^{2n+1}\times\mathbb{R}$ to $J^{1}\mathbb{R}^{2n+1}$ used in [Bhu01, San10, San11a] to the locally conformal symplectizations to obtain the desired identification (see equation (11) in Section 5). By considering the image by this identification of the lcs graph of a lcs Hamiltonian diffeomorphism $\varphi$ of $S^{1}\times\mathbb{R}^{2n+1}$ we obtain a Lagrangian submanifold $L_{\varphi}$ of $T_{-d\theta}(S^{1}\times\mathbb{R}^{2n+1})$. If $\varphi$ is compactly supported then $L_{\varphi}$ can be compactified to a Lagrangian submanifold $\overline{L_{\varphi}}$ of $T_{-d\theta}(S^{1}\times S^{2n+1})$, lcs Hamiltonian isotopic to the zero section. Similarly, we can associate to a compactly supported lcs Hamiltonian diffeomorphism $\varphi$ of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$ a Lagrangian submanifold $L_{\varphi}$ of $T_{-d\theta}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$, which can be compactified to a Lagrangian submanifold $\overline{L_{\varphi}}$ of $T_{-d\theta}(S^{1}\times S^{2n}\times S^{1})$, lcs Hamiltonian isotopic to the zero section. Using generating functions quadratic at infinity for $\overline{L_{\varphi}}$ we define, exactly as in the contact case for contactomorphisms of $\mathbb{R}^{2n+1}$ and $\mathbb{R}^{2n}\times S^{1}$ [Bhu01, San11a, San10] spectral selectors $c_{+}(\varphi)\geq 0$ and $c_{-}(\varphi)\leq 0$ for any compactly supported lcs Hamiltonian diffeomorphism $\varphi$ of $S^{1}\times\mathbb{R}^{2n+1}$ or $S^{1}\times\mathbb{R}^{2n}\times S^{1}$. These spectral selectors have the same properties as in the contact case (Proposition 6.3), and thus can be used to obtain similar applications. To start with, spectrality ($c_{+}(\varphi)$ and $c_{-}(\varphi)$ are equal to the action of an essential translated point of $\varphi$) and non-degeneracy ($c_{+}(\varphi)=c_{-}(\varphi)=0$ if and only if $\varphi$ is the identity) imply Theorem 1.2 (ii). Furthermore, as we will see, the properties listed in Proposition 6.3 imply that the partial relation $\preceq$ on the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n+1}$ or $S^{1}\times\mathbb{R}^{2n}\times S^{1}$ defined by posing $\varphi_{1}\preceq\varphi_{2}$ if $c_{+}(\varphi_{1}\circ\varphi_{2}^{-1})=0$ is a bi-invariant partial order (Theorem 7.1). By comparing the partial relation $\preceq$ with the one $\leq$ used to define orderability we then obtain a proof of Theorem 1.3 (ii). Moreover, we show that the arguments in [San11a, San10] can be generalized to the lcs case to obtain an unbounded integer-valued bi-invariant metric $d$ on the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, compatible with the partial orders $\preceq$ and $\leq$, by posing
@@
(Theorems 7.3 and 7.4). A lcs analogue of a known argument in contact geometry shows that the group of compactly supported lcs diffeomorphisms of any lcs manifold $\big(M,[(\eta,\omega)]\big)$ does not admit any bi-invariant metric that takes values arbitrarily close to zero (See Remark 7.5). This argument, however, does not directly apply to the group of lcs Hamiltonian diffeomorphisms. This motivates the following question.
Question. Does there exist a bi-invariant metric on $\operatorname{Ham}(S^{1}\times\mathbb{R}^{2n}\times S^{1})$ that takes values arbitrarily close to zero?
Finally, we show that the arguments in [San11a, San10] can be generalized to the lcs case to obtain an integer-valued lcs capacity for domains, by posing
@@
This will be the object of a forthcoming work. The above discussion also motivates the following question.
Question. Does there exist a lcs diffeomorphism $\varphi$ of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$ such that $\varphi(\overline{S^{1}\times B^{2n}(R_{1})\times S^{1}})\subset S^{1}\times B^{2n}(R_{2})\times S^{1}$ with $1\leq R_{2}<R_{1}$?
When the lcs manifold is $\big(\mathbb{R}^{2n},[(0,\omega_{0})]\big)$ then such a lcs diffeomorphism is provided by the Liouville flow. However, we currently do not know whether such a lcs squeezing exists for $S^{1}\times\mathbb{R}^{2n}\times S^{1}$.
@@
The article is organized as follows. In Section 2 we gather the basic notions on lcs manifolds that are needed in the rest of the article. Moreover, we discuss in 2.11 the notion of orderability and prove Theorem 1.3 (i). In Section 3 we define essential Lee chords and translated points. In Section 4 we recall from [CM19] the theory of generating functions for Lagrangian submanifolds of twisted cotangent bundles, and prove Theorem 1.1 and Theorem 1.2 (i). In Section 5 we define generating functions for compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n+1}$ and $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, and in Section 6 we use them to define the spectral selectors $c_{+}$ and $c_{-}$. A first application of these spectral selectors, Theorem 1.2 (ii), is also discussed in Section 6. In Section 7 we use the spectral selectors to define a partial order on the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n+1}$ and $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, and a bi-invariant metric on the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, and obtain a proof of Theorem 1.3 (ii). In Section 8 we define a lcs capacity for domains of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$, and use it to prove the lcs non-squeezing theorem for integers (Theorem 1.4). Finally, in Section 9 we discuss, following [SS24], how this non-squeezing result implies that the group of compactly supported lcs Hamiltonian diffeomorphisms of $S^{1}\times\mathbb{R}^{2n}\times S^{1}$ does not have the Rokhlin property.